Suppose the following economy produced pencils and erasers in 2011 and 2012 according to the following table. CPl basket in this economy consists of 5 pencils and 7 erasers. Which one is the value nearest to an inflation rate by CPI between 2010 and 2011 using 2010 as a base year? Q2010 P2010 02011 P2011 Pencils 5 10 8 30 Erasers 7 20 15 40 Select one: a. 135 O b. 100 O c. 115 d. 125
Answer #1

In 2010,

Price of basket of goods (5 pencils and 7 erasers) = 5*10 + 7*20 = 190

In 2011,

Price of the same basket = 5*30 + 7 *40 = 430

CP11 Inflation Rate-CP CPIT

430-190-126% Inflation Rate- 190

Nearest answer is Option d (125)